How many years of periodic, equal payments of $1,000 would be required to be made into an account earing 9% compounded annually, at the end of each compounding period to have on deposit $29,361 to pay future debt of the business?
1
Expert's answer
2019-09-24T11:34:36-0400

Term of an ordinary annuity is

n = ln(1 + FV×r/P)/ln(1 + r) = ln(1 + 29,361×0.09/1,000)/ln(1 + 0.09) = 15.